The revenue function​ R(x) and the cost function​ C(x) for a particular product are given. These functions are valid only for the specified range of values. Find the number of units that must be produced to break even.
R(x)=200x-x^2; C(x)=75x+2100; 0 less than or equal to x less than or equal to 100.
The manufacturer must produce ___ units to break even.

Answer by mananth(16946)


R(x)=200x-x^2; C(x)=75x+2100;
(0 less than or equal to x less than or equal to 100.)
for breakeven R(x)=C(x)
200x-x^2=75x+2100
simplify
x^2-175x +3100 =0
(x-20)(x-155)=0
x=20 or 155
but x cannot be grater than 100
so x=20